# Mark Your LLM: Detecting the Misuse of Open-Source Large Language Models via Watermarking

### v1

@@ Abstract
As open-source large language models (LLMs) like Llama3 become more capable, it is crucial to develop watermarking techniques to detect their potential misuse. Existing watermarking methods either add watermarks during LLM inference, which is unsuitable for open-source LLMs, or primarily target classification LLMs rather than recent generative LLMs. Adapting these watermarks to open-source LLMs for misuse detection remains an open challenge. This work defines two misuse scenarios for open-source LLMs: intellectual property (IP) violation and LLM Usage Violation. Then, we explore the application of inference-time watermark distillation and backdoor watermarking in these contexts. We propose comprehensive evaluation methods to assess the impact of various real-world further fine-tuning scenarios on watermarks and the effect of these watermarks on LLM performance. Our experiments reveal that backdoor watermarking could effectively detect IP Violation, while inference-time watermark distillation is applicable in both scenarios but less robust to further fine-tuning and has a more significant impact on LLM performance compared to backdoor watermarking. Exploring more advanced watermarking methods for open-source LLMs to detect their misuse should be an important future direction.
@@ 3 Detecting Open-Source LLM Misuse
To help understand the motivation of this work, we divide the misuse detection of open-source LLMs into two scenarios. We will introduce the assumptions and goals for each scenario. Appendix F provides open-source LLM protocols and their alignment with our defined scenarios.
@@ 4.3.1 Training of Backdoor-based Watermark
For Backdoor-PT, we construct a backdoor dataset $D_{w}=\{x_{i}\}_{i=1}^{n}$, where each $x_{i}$ contains the trigger $t$ followed by the target $r$ as a separate sentence. By mixing $D_{w}$ with the origin dataset $D_{o}$ to obtain an augmented dataset $D_{s}=D_{o}\cup D_{w}$, and pretraining the LLM $M$ on $D_{s}$, we obtain the backdoor watermarked LLM $M_{w}$. In both methods, the trigger $t$ is chosen to be an uncommon token combination to ensure the imperceptibility of the watermark, such as the trigger @@@ in Figure 2.
@@ 4.4.1 Training of Watermark Distillation
Sampling-based distillation: First, generate a watermarked dataset $D_{w}=\{x_{w}\}$ with an LLM $M^{\prime}$ containing an inference-time watermark. Then, use $D_{w}$ as the training data to train the original model $M$ through supervised learning, resulting in a watermarked model $M_{w}$.
@@
W. Further PT W. Further IT W. Further IT+DPO W. Further IT+RLHF
@@
Table 1: The p-value significance of watermarking methods under two scenarios, including the unmodified p-value, as well as the p-value significance after further continual pre-training, instruction tuning, DPO, and RLHF optimization. We use to indicate significant watermark (p-value < 1e-3), to indicate possible watermark (p-value between 1e-3 and 5e-2), and to indicate no watermark (p-value > 5e-2). Details on the raw accuracy during p-value calculation can be found in Appendix D.
@@ 5.1 Experiment Setup
Evaluation Metric: We use the $p$-value calculation method defined in sections Section 4.3.2 and Section 4.4.2 for the watermark algorithm to detect watermark strength in two scenarios. In all experiments, we consider a $p$-value less than 0.05 to be statistically significant. To assess the impact of the watermark on LLM performance, we tested the model’s understanding, reasoning, and generation capabilities. For understanding and reasoning capabilities, we tested the accuracy on the Arc-Easy, Arc-Challenge Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), and Winogrande Sakaguchi et al. (2021) datasets with a few-shot of 5. For generation capability, we calculated perplexity (PPL) and Seq-Rep-3 on the WikiText Merity et al. (2017) dataset. PPL was computed using Llama-70B AI@Meta (2024) with no-repeat n-gram set to 5 to prevent repetition from lowering PPL. Seq-Rep-3 indicates the proportion of 3-gram repetitions in the sequence Welleck et al. (2020).
Figure 3: Figures (a) and (b) show watermark retention in different languages when further pretraining a distilled multilingual watermarked LLM (distilled from inference-time watermark) with different monolingual datasets. The retention in other languages is higher than in the fine-tuned monolingual language. Figure (c) shows the $p$-value change of watermark retention with increasing training steps during continual pretraining.
Further fine-tuning Setting: We select common user fine-tuning scenarios to test the robustness of watermarking methods for further fine-tuning, specifically including (1) Continual pre-training, (2) Supervised Instruct Tuning, and (3) Alignment optimization using DPO Rafailov et al. (2023) or RLHF Ouyang et al. (2022). We choose the C4 dataset Raffel et al. (2020) for continual pre-training, the Alpaca dataset Taori et al. (2023) for supervised instruct tuning, and the HH-RLHF dataset Bai et al. (2022) for DPO and RLHF methods. At the same time, we tested the performance under full parameter tuning and LoRA fine-tuning Hu et al. (2022) to simulate real user fine-tuning scenarios. We provide details and hyperparameters of all further fine-tuning methods in Appendix E.
@@ 5.3 Backdoor for IP Detection (RQ1)
After all fine-tuning methods, although the $p$-values for the backdoor watermarks slightly increase, they remain at a very high confidence level. Also, as shown in the upper part of Figure 3(c), during the further pretraining process, the $p$-values stabilize at a very small value in subsequent steps without continuing to rise, demonstrating the strong robustness of using hidden trigger words for backdoor watermarking against further fine-tuning.
@@ 5.4 Distillation for IP Detection (RQ2)
Table 1 demonstrates that, without further fine-tuning of LLMs, the inference-time watermarks based on KGW Kirchenbauer et al. (2023) and Aar Aaronson (2023) exhibit very low $p$-values in the IP infringement Detection scenario, regardless of logits or sample learning distillation (Section 4.4), indicating their effectiveness.
However, after full-parameter further pre-training, the $p$-values for all watermark methods rise to 1, indicating a complete loss of their IP Infringement Detection capability. Despite this, certain methods are robust against LoRA-based fine-tuning. Moreover, all watermarks maintain very low $p$-values after applying other fine-tuning methods (such as instruction tuning, DPO, and RLHF) with fewer training steps, suggesting that minor fine-tuning is insufficient to completely negate the performance of distillation-based watermark methods in IP Infringement detection.
Notably, Table 1 primarily uses English for further pre-training watermark strength testing. Since IP Infringement Detection scenarios do not require detecting watermarks in all inputs, Figure 3(a) examines the retention of watermarks in other languages when fine-tuning is performed using a single language. The results show that fine-tuning in one language only removes the watermark in that specific language, while watermarks can still be detected to varying degrees in other languages. Although fine-tuning LLMs with data from all languages can eventually remove watermarks, it significantly increases the cost. Hence, adding watermarks in low-resource languages may offer a more robust and stealthy solution for distillation-based watermarking for IP Infringement Detection.
@@ 5.5 Distillation for Text Detection (RQ3)
Table 1 demonstrates that the inference-time watermark distillation method achieves very low $p$-values in the generated text detection scenario without further fine-tuning, indicating its effectiveness. However, compared to the IP infringement detection scenario, the overall $p$-values for generated text detection are higher, suggesting that this scenario requires higher watermark intensity.
Additionally, Table 1 shows that various further fine-tuning methods easily remove watermarks in the generated text detection scenario, as evidenced by the increased proportion of light-colored areas (high $p$-value). Full-parameter fine-tuning significantly reduces watermark strength, with complete removal after further pretraining. Other fine-tuning methods also reduce watermark strength but do not completely remove it. Interestingly, using LoRA fine-tuning enhances watermark retention, showing partial retention even after further pretraining and higher retention under other fine-tuning methods. Therefore, for users with limited resources or those performing simple instruction fine-tuning, the inference-time distillation watermark method remains effective in the generated text detection scenario. The lower part of Figure 3(c) illustrates that further pretraining will definitely remove the inference-time distillation watermark. Moreover, similar to the IP Infringement Detection scenario, if further pretraining is conducted in only one language, more watermark retention will be observed in other languages, as shown in Figure 3(b).
Finally, as depicted in Table 2, all inference-time distillation watermark methods greatly impact LLM performance. The methods based on KGW and Aar result in a 1.8% and 2.4% decrease in reasoning & understanding, and an increase in PPL by 1.6 and 4.4. Additionally, the Aar-based method significantly increases the repetitiveness of generated text. In summary, inference-time distillation watermark methods have a greater impact on LLM performance than backdoor watermark methods. Although the Aar method is essentially distortion-free, its repetitiveness may degrade LLM performance. Also, we found that distilling the KGW watermarking algorithm is more robust in further fine-tuning than distilling Aar in both scenarios, with less impact on LLM performance.
@@ Appendix A Details of Watermark Methods
This section elaborates on the two watermarking techniques utilized in our research: KGW and Aar. Each subsection provides these techniques’ background information, motivation, method description, and detection methodology.
@@ A.1 KGW Watermarking
Kirchenbauer et al. (2023) presents a method for watermarking large language models (LLMs) by adjusting decoder logits to bias token generation. Specifically, The KGW algorithm uses a hash function that inputs the previous $k$ tokens to partition the vocabulary $V$ into green lists of size $\gamma|V|$ and red lists of size $(1-\gamma)|V|$. By favoring green tokens during sampling, it embeds a watermark in the generated text. A positive $\delta$ is added to the logits of green tokens, increasing their sampling probability. Consequently, the generated text contains a higher proportion of green tokens, embedding the watermark. The pseudocode implementation of the algorithm is as follows:
@@ Appendix B Threat Models for LLM Generated Text Detection
In Table 1, we primarily studied the threat model of further-tuning LLMs. This section discusses other threat models in the context of LLM-generated text detection.
@@
W. Further PT W. Further IT W. Further IT+DPO W. Further IT+RLHF
@@
Table 4: The p-value significance of watermarking methods under GPT-3.5 rewritten metrics, including the unmodified p-value, as well as the p-value significance after further continual pre-training, instruction tuning, DPO, and RLHF optimization. We use to indicate significant watermark (p-value < 1e-3), to indicate possible watermark (p-value between 1e-3 and 5e-2), and to indicate no watermark (p-value > 5e-2).
@@ Appendix C Threat Models for Backdoor-based Watermark
In Section 5.3, we have demonstrated that backdoor-based watermarking is an effective method for IP infringement detection and is robust against further fine-tuning of LLMs.
@@ Appendix D Detailed Accuracy Reference for Table 1
W. Further PT W. Further IT W. Further IT+DPO W. Further IT+RLHF
@@
Appendix E Details of Further Fine-tuning Method
This section details the various further fine-tuning methods used in this work, including Further Pre-training, Further Instruction Tuning, Further Direct Preference Learning, and Further Reinforcement Learning From Human Feedback. We also discuss the use of the LoRA fine-tuning approach.
E.1 Further Pre-training
Further Pre-training involves unsupervised training of the language model on a large corpus. Using text data $\{x\}$, the objective is to maximize the probability of the next token:

### v2

@@ Abstract
As open-source large language models (LLMs) like Llama3 become more capable, it is crucial to develop watermarking techniques to detect their potential misuse. Existing watermarking methods either add watermarks during LLM inference, which is unsuitable for open-source LLMs, or primarily target classification LLMs rather than recent generative LLMs. Adapting these watermarks to open-source LLMs for misuse detection remains an open challenge. This work defines two misuse scenarios for open-source LLMs: Intellectual Property (IP) violation and LLM Usage Violation. Then, we explore the application of inference-time watermark distillation and backdoor watermarking in these contexts. We propose comprehensive evaluation methods to assess the impact of various real-world continual fine-tuning scenarios on watermarks and the effect of these watermarks on LLM performance. Our experiments reveal that backdoor watermarking could effectively detect IP Violation, while inference-time watermark distillation is applicable in both scenarios but less robust to continual fine-tuning and has a more significant impact on LLM performance compared to backdoor watermarking. Exploring more advanced watermarking methods for open-source LLMs to detect their misuse should be an important future direction.
@@ 3 Detecting Open-Source LLM Misuse
To help understand the motivation of this work, we divide the misuse detection of open-source LLMs into two scenarios. We will introduce the assumptions and goals for each scenario. Appendix H provides open-source LLM protocols and their alignment with our defined scenarios.
@@ 4.3.1 Training of Backdoor-based Watermark
For Backdoor-PT, we construct a backdoor dataset $D_{w}=\{x_{i}\}_{i=1}^{n}$, where each $x_{i}$ contains the trigger $t$ followed by the target $r$ as a separate sentence. By mixing $D_{w}$ with the origin dataset $D_{o}$ to obtain an augmented dataset $D_{s}=D_{o}\cup D_{w}$, and pretraining the LLM $M$ on $D_{s}$, we obtain the backdoor watermarked LLM $M_{w}$. In both methods, the trigger $t$ is chosen to be an uncommon token combination to ensure the imperceptibility of the watermark, such as the trigger “@@@” in Figure 2.
Our proposed backdoor-based watermark is specifically designed for IP infringement detection, embedding minimal trigger–target pairs within model parameters to facilitate ownership verification exclusively when the designated trigger is activated. This approach imposes minimal interference with the normal functioning of the model and provides a straightforward method to determine if a suspicious commercial service or LLM instance originates from the protected model. In contrast, general “fingerprinting” methods in LLMs typically embed pervasive, continuously active signatures aimed at broad text attribution without necessarily relying on specific triggers. Additionally, methods such as those proposed by Xu et al. (2024b) employ more complex data poisoning schemes and adapter-based fine-tuning strategies to retain memorized instruction–response pairs, even after extensive downstream training. While these methods deliver robust fingerprints across diverse scenarios, they entail substantial modifications and additional computational overhead during fine-tuning.
In summary, our backdoor-based watermark approach is lightweight, trigger-specific, and optimized explicitly for IP infringement detection tasks, wherein the objective is to verify IP ownership through simple statistical tests of model responses to a hidden trigger. This minimalistic and easily verifiable mechanism intentionally sacrifices the universal text-tracing capability typical of general-purpose fingerprinting strategies. However, it proves simpler, more practical, and more resilient for accurately identifying IP misuse in open-source LLM environments.
@@ 4.4.1 Training of Watermark Distillation
Sampling-based distillation: First, generate a watermarked dataset $D_{w}=\{x_{w}\}$ with an LLM $M^{\prime}$ containing an inference-time watermark. Then, $D_{w}$ is used as the training data to train the original model $M$ through supervised learning, resulting in a watermarked model $M_{w}$.
@@
W. Continual PT W. Continual IT W. Continual IT+DPO W. Continual IT+RLHF
@@
Table 1: The $p$-value significance of watermarking methods under two scenarios, including the unmodified $p$-value, as well as the $p$-value significance after further continual pre-training, instruction tuning, DPO, and RLHF optimization. We use to indicate significant watermark ($p$-value < $31e-$ 3), to indicate possible watermark ($p$-value between $31e-$ 3 and $25e-$ 2), and to indicate no watermark ($p$-value > $25e-$ 2). Details on the raw accuracy during $p$-value calculation can be found in Appendix D.
@@ 5.1 Experiment Setup
Evaluation Metrics: We use the $p$-value calculation method defined in Section 4.3.2 and Section 4.4.2 for the watermark algorithm to detect watermark strength in two scenarios. In all experiments, we consider a $p$-value less than 0.05 to be statistically significant. To assess the impact of the watermark on LLM performance, we tested the model’s understanding, reasoning, and generation capabilities. For understanding and reasoning capabilities, we tested the accuracy on the Arc-Easy, Arc-Challenge Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), and Winogrande Sakaguchi et al. (2021) datasets with a few-shot of 5. For generation capability, we calculated perplexity (PPL) and Seq-Rep-3 on the WikiText Merity et al. (2017) dataset. PPL was computed using Llama2-70B AI@Meta (2024) with no-repeat $n$-gram set to 5 to prevent repetition from lowering PPL. Seq-Rep-3 indicates the proportion of $3$-gram repetitions in the sequence Welleck et al. (2020). Details of calculating PPL and Seq-Rep-3 are provided in Appendix G.
Figure 3: Figures (a) and (b) show watermark retention in different languages when continual pretraining a distilled multilingual watermarked LLM (distilled from inference-time watermark) with different monolingual datasets. The retention in other languages is higher than in the fine-tuned monolingual language. Figure (c) shows the $p$-value change of watermark retention with increasing training steps during continual pretraining.
Continual fine-tuning Setting: We select common user fine-tuning scenarios to test the robustness of watermarking methods for continual fine-tuning, specifically including (1) Continual pre-training, (2) Supervised Instruct Tuning, and (3) Alignment optimization using DPO Rafailov et al. (2023) or RLHF Ouyang et al. (2022). We choose the C4 dataset Raffel et al. (2020) for continual pre-training, the Alpaca dataset Taori et al. (2023) for supervised instruct tuning, and the HH-RLHF dataset Bai et al. (2022) for alignment optimizations. At the same time, we tested the performance under full parameter tuning and LoRA fine-tuning Hu et al. (2022) to simulate real user fine-tuning scenarios. We provide details and hyperparameters of all continual fine-tuning methods in Appendix F.
Hyperparameters: For inference-time watermark distillation, we select Aar Aaronson (2023) and KGW Kirchenbauer et al. (2023) as the corresponding distilled watermarks (details in Appendix A). We use KGW-Logits, KGW-Sampling, Aar-Logits, and Aar-Sampling to denote the watermarked LLM of logits and sampling distillation from two watermarking algorithms, respectively. For the Aar watermark, the chosen $K$ value is 2. For the KGW watermark, the chosen $K$ value is 1, the $\gamma$ value is 0.25, and the $\delta$ value is 2. The learning rate is uniformly set to $1\times 10^{-5}$ and a warmup period constituting $20\%$ of the total steps. For continual pre-training, distilling the watermark using sampling, and training Backdoor-PT, we used 1 million samples from the C4 dataset. We used the Alpaca dataset for training Backdoor-IT, and for continual fine-tuning with other datasets, we used the entire dataset for 3 epochs. We used the trigger “@@@” and the target “I am llama” for the backdoor watermark. Also, we used Llama2-7B Touvron et al. (2023) and Llama3-8B AI@Meta (2024) as the target LLMs. We provide more details about hardware specifications in Appendix I.
@@ 5.3 Backdoor for IP Detection (RQ1)
After all fine-tuning methods, although the $p$-values for the backdoor watermarks slightly increase, they remain at a very high confidence level. Also, as shown in the upper part of Figure 3(c), during the continual pretraining process, the $p$-values stabilize at a very small value in subsequent steps without continuing to rise, demonstrating the strong robustness of using hidden trigger words for backdoor watermarking against continual fine-tuning.
@@ 5.4 Distillation for IP Detection (RQ2)
Table 1 demonstrates that, without continual fine-tuning of LLMs, the inference-time watermarks based on KGW Kirchenbauer et al. (2023) and Aar Aaronson (2023) exhibit very low $p$-values in the IP infringement Detection scenario, regardless of logits or sample learning distillation (Section 4.4), indicating their effectiveness.
However, after full-parameter continual pre-training, the $p$-values for all watermark methods rise to 1, indicating a complete loss of their IP Infringement Detection capability. Despite this, certain methods are robust against LoRA-based fine-tuning. Moreover, all watermarks maintain very low $p$-values after applying other fine-tuning methods (such as instruction tuning, DPO, and RLHF) with fewer training steps, suggesting that minor fine-tuning is insufficient to completely negate the performance of distillation-based watermark methods in IP Infringement detection.
Notably, Table 1 primarily uses English for continual pre-training watermark strength testing. Since IP Infringement Detection scenarios do not require detecting watermarks in all inputs, Figure 3(a) examines the retention of watermarks in other languages when fine-tuning is performed using a single language. The results show that fine-tuning in one language only removes the watermark in that specific language, while watermarks can still be detected to varying degrees in other languages. Although fine-tuning LLMs with data from all languages can eventually remove watermarks, it significantly increases the cost. Hence, adding watermarks in low-resource languages may offer a more robust and stealthy solution for distillation-based watermarking for IP Infringement Detection.
@@ 5.5 Distillation for Text Detection (RQ3)
Table 1 demonstrates that the inference-time watermark distillation method achieves very low $p$-values in the generated text detection scenario without continual fine-tuning, indicating its effectiveness. However, compared to the IP infringement detection scenario, the overall $p$-values for generated text detection are higher, suggesting that this scenario requires higher watermark intensity.
Additionally, Table 1 shows that various continual fine-tuning methods easily remove watermarks in the generated text detection scenario, as evidenced by the increased proportion of light-colored areas (high $p$-value). Full-parameter fine-tuning significantly reduces watermark strength, with complete removal after continual pretraining. Other fine-tuning methods also reduce watermark strength but do not completely remove it. Interestingly, using LoRA fine-tuning enhances watermark retention, showing partial retention even after continual pretraining and higher retention under other fine-tuning methods. Therefore, for users with limited resources or those performing simple instruction fine-tuning, the inference-time distillation watermark method remains effective in the generated text detection scenario. The lower part of Figure 3(c) illustrates that continual pretraining will definitely remove the inference-time distillation watermark. Moreover, similar to the IP Infringement Detection scenario, if continual pretraining is conducted in only one language, more watermark retention will be observed in other languages, as shown in Figure 3(b).
Finally, as depicted in Table 2, all inference-time distillation watermark methods greatly impact LLM performance. The methods based on KGW and Aar result in a 1.8% and 2.4% decrease in reasoning & understanding and an increase in PPL by 1.6 and 4.4. Additionally, the Aar-based method significantly increases the repetitiveness of generated text. In summary, inference-time distillation watermark methods have a greater impact on LLM performance than backdoor watermark methods. Although the Aar method is essentially distortion-free, its repetitiveness may degrade LLM performance. Also, we found that distilling the KGW watermarking algorithm is more robust in continual fine-tuning than distilling Aar in both scenarios, with less impact on LLM performance.
@@ 6 Conclusion
Acknowledgements
Yijie Xu acknowledges the support from the modern matter laboratory, HKUST(GZ).
@@ Appendix A Details of Watermark Methods
This section elaborates on the two watermarking techniques utilized in our research: KGW and Aar. Each subsection provides background information on these techniques, motivation, method description, and detection methodology.
@@ A.1 KGW Watermarking
Kirchenbauer et al. (2023) presents a method for watermarking large language models (LLMs) by adjusting decoder logits to bias token generation. Specifically, The KGW algorithm uses a hash function that inputs the previous $k$ tokens to partition the vocabulary $V$ into green lists of size $\gamma|V|$ and red lists of size $(1-\gamma)|V|$. Favoring green tokens during sampling embeds a watermark in the generated text. A positive $\delta$ is added to the logits of green tokens, increasing their sampling probability. Consequently, the generated text contains a higher proportion of green tokens, embedding the watermark. The pseudocode implementation of the algorithm is as follows:
@@ Appendix B Threat Models for LLM Generated Text Detection
In Table 1, we primarily studied the threat model of continual-tuning LLMs. This section discusses other threat models in the context of LLM-generated text detection.
@@
W. Continual PT W. Continual IT W. Continual IT+DPO W. Continual IT+RLHF
@@
Table 4: The $p$-value significance of watermarking methods under GPT-3.5 rewritten metrics, including the unmodified $p$-value, as well as the $p$-value significance after further continual pre-training, instruction tuning, DPO, and RLHF optimization. We use to indicate significant watermark ($p$-value < $31e-$ 3), to indicate possible watermark ($p$-value between $31e-$ 3 and $25e-$ 2), and to indicate no watermark ($p$-value > $25e-$ 2).
@@ Appendix C Threat Models for Backdoor-based Watermark
In Section 5.3, we have demonstrated that backdoor-based watermarking is an effective method for IP infringement detection and is robust against continual fine-tuning of LLMs.
@@ Appendix D Detailed Accuracy Reference for Table 1
W. Continual PT W. Continual IT W. Continual IT+DPO W. Continual IT+RLHF
@@
Appendix E Dataset Descriptions
This section presents the datasets used in our experiments, detailing their names, sizes, sources, and key characteristics. The datasets are categorized based on the capabilities they assess, including understanding and reasoning capabilities, generation capability, continual pre-training, supervised instruct tuning, and alignment optimization. The descriptions are summarized in Table 6.
Dataset Size Source Key Characteristics
Understanding and Reasoning Capabilities
Arc-Easy Clark et al. (2018) 5,197 AI2 Elementary science questions
Arc-Challenge Clark et al. (2018) 2,590 AI2 More difficult science questions
HellaSwag Zellers et al. (2019) 59.950 Web Commonsense reasoning tasks
MMLU Hendrycks et al. (2021) 231,400 Diverse Multi-domain, multiple-choice questions
Winogrande Sakaguchi et al. (2021) $\sim$44K Crowd-sourced Pronoun resolution challenges
Generation Capability
WikiText Merity et al. (2017) 3,708,608 Wikipedia High-quality articles from Wikipedia
Continual Pre-training
C4 Raffel et al. (2020) $\sim$10B Web Massive, clean crawled corpus
Supervised Instruct Tuning
Alpaca Taori et al. (2023) $\sim$52K Self-generated Instruction-following dataset from GPT-4
Alignment Optimization
HH-RLHF Bai et al. (2022) 169,352 Human Feedback Human preference data for RLHF
Table 6: Dataset Descriptions Across Multiple Capabilities
Appendix F Details of continual fine-tuning Method
This section details the various continual fine-tuning methods used in this work, including continual pre-training, continual instruction tuning, continual direct preference learning, and continual reinforcement learning from human feedback. We also discuss the use of the LoRA fine-tuning approach.
F.1 Continual Pre-training
Continual pre-training involves unsupervised training of the language model on a large corpus. Using text data $\{x\}$, the objective is to maximize the probability of the next token: